hu 18 synthesis of glycine from ammonia. For glycine, 537.2 kJ mol Assume T 298 K. e. Calculate the standard enthalpy change for thraynthesis of glycine from nitrogen, water, oxygen, and methane If 3.365 g ethanol CHHsoHa is burned com- pletely in a bomb calorimeter at 298.15 K. the heat produced rest, is 99,472 kl. for ethanol at 298.15 K calculate AH reac. b. Calculate AHT of ethanol at 298.15 K. (g) is P4.21 From the following data, calculate AHR AK for the reaction CHrcooHkg) t 2020e follow. Hzou values for sHh for the first two reactions are at 2gs 15 K. and for the third reaction at 391,4 K substance CH cool) one) covey Haob Haoe P4.22 Ati sample o


4.20
Answer #1

a)

Hcombustion = Qcombustion / moles

Q = 99.472 kJ

n = mass/MW = 3.365/46 = 0.0731521 mol of ethanol

HRcomb = (99.472/0.0731521) = 1359.79 kJ/mol

b)

Hf for ehtanol given:

C2H5OH(l) + O2 (g)= CO2(g) + H2O(g)

balance:

C2H5OH(l) + 3O2 (g)= 2CO2(g) + 3H2O(g)

HRxn = Hproducts - Hreactants

1359.79 = (2*-393.5 + 3*-241.818) - (C2H5OH + 3*0)

C2H5OH = 1359.79- (2*-393.5 + 3*-241.818)

C2H5OH = 2,872.24 kJ/mol